# Communication-Efficient and Privacy-Adaptable Mechanism for Federated Learning

### v1

@@ V-A4 Training Configurations
V-B FL Convergence
We report the FL convergence of CEPAM-Gaussian and CEPAM-Laplace in terms of accuracy using MLP and CNN architectures and MNIST dataset.
V-B1 Gaussian Mechanism
We set $\sigma=0.1$ and $\tilde{\epsilon}=5.9$ for the base Gaussian mechanisms in CEPAM-Gaussian for $n=1,2,3$ with variances $0.1^{2}$, $2\times 0.1^{2}$, $3\times 0.1^{2}$, respectively, and all the clients’ composite Gaussian mechanism achieve $(\epsilon=1.45,\delta=9.69\times 10^{-3})$-DP. We also set $\sigma=0.1$ for both FL+Gaussian and FL+Gaussian+SDQ.
@@ V-B1 Gaussian Mechanism
(a) Convergence profiles of different FL schemes using MLP.
(b) Convergence profiles of different FL schemes using CNN.
Figure 1: Validation accuracy curves of CEPAM-Gaussian against Baselines using different neural network models.
In Table I, we report the test accuracy of CEPAM-Gaussian for different architectures (MLP and CNN) using MNIST. CEPAM-Gaussian demonstrates better performance, achieving an improvement of 1.0-2.0% and 0.5-1.0% in accuracy compared to other baselines for MLP and CNN, respectively, suggesting that CEPAM can be beneficial for various learning models. Furthermore, all cases of CEPAM-Gaussian have approximately the same accuracy. This is because the added noise per dimension is consistent, i.e., $\mathrm{Var}(f)/n=\sigma^{2}$. However, higher-dimensional LRSUQ has a better compression ratio than lower-dimensional counterparts due to its nature as a vector quantizer. In contrast, for baselines such as the Gaussian-mechanism-then-quantize approach (Gaussian+SDQ), the total noise comprises privacy noise and quantization error, leading to more distortion and consequently lower accuracy compared to LRSUQ. It is worth noting that adding a minor level of distortion as a regularizer during training deep models can potentially enhance the performance of the converged model [63]. This finding is corroborated in our experimental results, where CEPAM-Gaussian achieves slightly better accuracy than vanilla FL for both architectures.
Table I: Test Accuracy (Gaussian) for MNIST
@@
FL 82.50 91.34
FL+SDQ 82.42 91.46
FL+Gaussian 83.14 91.39
FL+Gaussian+SDQ 83.15 90.74
CEPAM-Gaussian ($n=1$) 84.42 91.83
CEPAM-Gaussian ($n=2$) 84.53 91.61
CEPAM-Gaussian ($n=3$) 83.85 91.94
V-B2 Laplace Mechanism
We set $b=0.1$ and $\tilde{\epsilon}=5000$ for the base Laplace mechanism in CEPAM-Laplace with a variance of $2\times 0.1^{2}$, and the composite Laplace mechanism achieves $(\epsilon=4995)$-DP. We also set $b=0.1$ for both FL+Laplace and FL+Laplace+SDQ. Note that a higher privacy budget is required to ensure $\epsilon$-DP for CEPAM-Laplace under privacy amplification with subsampling analysis. However, we might achieve a lower privacy budget by relaxing $\delta$. The trade-off is left for the future study.
Figure 2 shows the validation accuracy of CEPAM-Laplace over communication rounds using the MLP and CNN models, respectively. Similar to CEPAM-Gaussian, we observe that CEPAM-Laplace consistently outperforms other baselines.
(a) Convergence profiles of different FL schemes using MLP.
(b) Convergence profiles of different FL schemes using CNN.
Figure 2: Validation accuracy curves of CEPAM-Laplace against Baselines using different neural network models.
In Table II, we also report the performance of test accuracy of CEPAM-Laplace. We observe that CEPAM-Laplace achieves an improvement of about 2.0-2.7% and 1.5-2.0% in accuracy compared to other baselines for MLP and CNN, respectively, which suggests that CEPAM-Laplace can be beneficial for many learning models.
Table II: Test Accuracy (Laplace) for MNIST
Baselines MLP (%) CNN (%)
FL 83.02 91.08
FL+SDQ 83.28 91.00
FL+Laplace 83.78 90.59
FL+Laplace+SDQ 83.78 90.86
CEPAM-Laplace 85.76 92.66
V-C Privacy-Accuracy Trade-off
Figure 3 illustrates the trade-off between learning performance, measured in terms of test accuracy, and the privacy budget between CEPAM and the simple Gaussian-mechanism-then-quantize approach (Gaussian+SDQ) using a CNN architecture. Overall, CEPAM-Gaussian outperforms Gaussian+SDQ. The figure demonstrates that as more privacy budget is allocated, higher test accuracy can be achieved with CEPAM-Gaussian. However, there is a point of diminishing returns; once the privacy budget reaches a threshold (i.e., $\epsilon=3.0$), the increase in test accuracy by further increasing the privacy budget becomes limited.
@@ V-C Privacy-Accuracy Trade-off
VI Conclusion and Discussion
In this paper, we introduced CEPAM, which can be used to achieved communication efficiency and privacy protection simultaneously in FL framework. We also proved the FL convergence bound for CEPAM and provided a privacy amplication analysis for CEPAM-Gaussiang and CEPAM-Laplace. Furthermore, the experimental results confirmed our theoretical findings through a comparison of CEPAM with alternative methods using the MNIST dataset with MLP and CNN architectures. The experimental outcomes demonstrated that CEPAM-Gaussian and CEPAM-Laplace lead to a moderate enhancement in accuracy by 0.5-2.0% and 1.5-2.7%, respectively, when contrasted with other baseline approaches.

### v2

@@ VI-A4 Training Configurations
VI-A5 Repetition Strategy
For each set of parameters and baseline method, the training process was repeated 10 times with varying random seeds, and the average performance across these runs was reported.
VI-B FL Convergence
We report the FL convergence of CEPAM-Gaussian in terms of accuracy using MLP and CNN architectures and MNIST dataset.
VI-B1 Gaussian Mechanism
We set $\sigma=0.001$ and $\tilde{\epsilon}=5.9$ for the base Gaussian mechanisms in CEPAM-Gaussian for $n=1,2,3$ with variances $0.001^{2}$, $2\times 0.001^{2}$, $3\times 0.001^{2}$, respectively. By Theorem 12, all the clients’ composite Gaussian mechanisms achieve $(\epsilon=1.45,\delta=2.48\times 10^{-2})$-DP. We also set $\sigma=0.001$ for both FL+Gaussian and FL+Gaussian+SDQ.
@@ VI-B1 Gaussian Mechanism
Figure 1: Convergence profile of different FL schemes.
In Table I, we report the test accuracy of CEPAM-Gaussian for different architectures (MLP and CNN) using MNIST with their 95% confidence intervals. CEPAM-Gaussian demonstrates better performance, achieving an improvement of 0.6-2.0% and 0.4-1.0% in accuracy compared to other baselines for MLP and CNN, respectively, suggesting that CEPAM can be beneficial for various learning models. Furthermore, all cases of CEPAM-Gaussian have approximately the same accuracy. This is because the added noise per dimension is consistent, i.e., $\mathrm{Var}(f)/n=\sigma^{2}$. However, higher-dimensional LRSUQ has a better compression ratio than scalar counterparts [21, 22] due to its nature as a vector quantizer. In contrast, for baselines such as the Gaussian-mechanism-then-quantize approach (Gaussian+SDQ), the total noise comprises privacy noise and quantization error, leading to more distortion and consequently lower accuracy compared to LRSUQ. It is worth noting that adding a minor level of distortion as a regularizer during training deep models can potentially enhance the performance of the converged model [78]. This finding is corroborated in our experimental results, where CEPAM-Gaussian achieves slightly better accuracy than vanilla FL for both architectures.
@@
FL $82.06\pm 1.35$ $91.70\pm 0.66$
FL+SDQ $82.17\pm 1.31$ $91.73\pm 0.68$
FL+Gaussian $82.23\pm 1.23$ $91.84\pm 0.72$
FL+Gaussian+SDQ $82.26\pm 1.20$ $91.84\pm 0.70$
CEPAM-Gaussian ($n=1$) $83.99\pm 0.92$ $92.74\pm 0.70$
CEPAM-Gaussian ($n=2$) $83.52\pm 1.00$ $92.26\pm 0.64$
CEPAM-Gaussian ($n=3$) $82.84\pm 1.19$ $92.29\pm 0.72$
Table I: Test Accuracy for MNIST
VI-C Privacy-Accuracy Trade-off
We evaluate the privacy-accuracy trade-off using two different values of privacy relaxation $\delta$. In the following, the parameters are computed according to Theorem 12. For $\delta=0.01$, we conducted the experiments by varying privacy budget $\epsilon$ from 1 to 10, the corresponding $\sigma$ are $\{0.1375,0.0864,0.0602,0.0431,0.0307,0.0231,0.0141,0.0083,0.00364,0.000159\}$ respectively. Similarly, for $\delta=0.015$, we conducted the experiments by varying privacy budget $\epsilon$ from 1 to 5, the corresponding $\sigma$ are $\{0.0857,0.0445,0.0236,0.0101,0.0005\}$ respectively. Again, for each set of parameters, the experiments are simulated 10 times, and the results are averaged.
Figure 2 illustrates the trade-off between learning performance, measured in terms of test accuracy, and the privacy budget between CEPAM and the simple Gaussian-mechanism-then-quantize approach (Gaussian+SDQ) using a CNN architecture.
Overall, CEPAM-Gaussian outperforms Gaussian+SDQ. The figure demonstrates that as more privacy budget is allocated, higher test accuracy can be achieved with CEPAM-Gaussian. However, for each $\delta$, there is a point of diminishing returns; once the privacy budget reaches a threshold. In particular, at $\epsilon\approx 7$ and $\epsilon\approx 4$ for $\delta=0.01,0.015$ respectively, the increase in test accuracy by further increasing the privacy budget becomes limited.
@@ VI-C Privacy-Accuracy Trade-off
VII Conclusion and Discussion
In this paper, we introduced CEPAM, which can be used to achieved communication efficiency and privacy protection simultaneously in FL framework. We also proved the FL convergence bound for CEPAM and provided a privacy amplification analysis for CEPAM-Gaussian and CEPAM-Laplace. Furthermore, the experimental results confirmed our theoretical findings through a comparison of CEPAM with alternative methods using the MNIST dataset with MLP and CNN architectures. The experimental outcomes demonstrated that CEPAM-Gaussian and CEPAM-Laplace lead to a moderate enhancement in accuracy by 0.5-2.0% and 1.5-2.7%, respectively, when contrasted with other baseline approaches.
@@ VII Conclusion and Discussion
Appendix A Proofs for Section V
A-A Proof of Lemma 11
Since we are using LRSUQ for the encoding and decoding of the $j$-th sub-vector $\tilde{\mathbf{X}}_{t+\tau,j}^{k}$, where $j\in\mathcal{N}$, Proposition 5 implies that, regardless of the statistical model of $\tilde{\mathbf{X}}_{t+\tau,j}^{k}$, the quantization errors $\{\tilde{\mathbf{Z}}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ are iid (over $k$ and $j$) and follows the pdf $f$.
A-B Proof of Theorem 12
Our proof relies on the following fact about the mixture of uniform distributions [33].
Lemma 14.
If $\mathbf{Z}\,|\,\{\tilde{U}=u\}\sim\mathrm{Unif}(\sigma\sqrt{u}B^{n})$ and $\tilde{U}\sim\chi^{2}_{n+2}$, then $\mathbf{Z}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{n})$.
In other words, given the noise vector $\mathbf{Z}$, which follows a zero-mean multivariate Gaussian distribution with a covariance matrix $\sigma^{2}\mathbf{I}_{n}$, where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix, i.e., $\mathbf{Z}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{n})$, $\mathbf{Z}$ can be expressed as a mixture of uniform distributions over $n$-dimensional balls, with the latent variable $\tilde{U}$ following a chi-squared distribution with $n+2$ degrees of freedom. By setting $\tilde{U}_{t+\tau,j}^{k}\sim g=\chi_{n+2}^{2}$ for every $k$ and $\tilde{\mathbf{Z}}_{t+\tau,j}^{k}\;|\;\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}(\sigma\sqrt{u}B^{n})$ for $k$ and $j$, Lemma 11 and Lemma 14 imply that the resulting mechanism, i.e., Step 7 in Algorithm 3, is Gaussian.
We utilize the subsampling with replacement mechanism to model the $\tau$-iterations of local SGD at $k$-th client, [^11] According to Equation (3), each $k$-th client performs $\tau$ iterations of local SGD with uniformly sampled indices from the dataset $\mathcal{D}^{(k)}$ (i.e., $p_{j}=1/|\mathcal{D}^{(k)}|=1/n$, where $j\in\{1,\ldots,n\}$ and $n=|\mathcal{D}^{(k)}|$ in (12) or (14)), this process can be regarded as subsampling with replacement, which outputs a multiset sample of size $\tau$ following a multinomial distribution $\mathrm{Mult}(\tau;p_{1},\ldots,p_{n})$. instead of employing Poisson subsampling for mini-batching [39] within a single local iteration as in [21]. Furthermore, the resulting subsampled Gaussian mechanism is applied to sub-vectors of dimension $n$ rather than to individual components.
For the $j$-th sub-vector in $\{\mathbf{Y}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ where $k\in\mathcal{K}$, we have $\tilde{\mathbf{Z}}_{t+\tau,j}^{k}\,|\,\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}(\sigma\sqrt{u}B^{n})$. Since $\tilde{U}_{t+\tau,j}^{k}\sim\chi^{2}_{n+2}$, the noisy sub-vectors $\tilde{\mathbf{Z}}_{t+\tau,j}^{k}$ follows a Gaussian distribution $\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{n})$ by Lemma 14. Given that this is true for any client $k^{\prime}\neq k$, the quantization noise of one client is Gaussian distributed from the perspective of any other client. Considering the averaging operation at the central server, we obtain (15) by Lemma 11, since sum of independent Gaussian random variables is a Gaussian random variable with appropriate parameters.
For the $(\epsilon,\delta)$-DP guarantee, we apply Lemma 9 on every $j$-th sub-vector $\hat{\mathbf{X}}_{t+\tau,j}^{k}$, which is the output of the subsampled Gaussian mechanism. Since the server averages sum of $\hat{\mathbf{X}}_{t+\tau,j}^{k}$’s over $K$ clients, the effect of a single data point $\xi\in\mathcal{D}^{(k)}$ is at most $\frac{\tau\gamma}{K}$. [^12] Since there are $\tau$ iterations, any single data point $\xi\in\mathcal{D}^{(k)}$ can appear in the subsample at most $\tau$ times. Consequently, the $\ell_{2}$ sensitivity of $\frac{1}{K}\sum_{k\in\mathcal{K}}\mathbf{X}_{t+\tau,j}^{k}$ is at most $\frac{2\tau\gamma}{K}$. Substituting $\Delta_{2}=\frac{2\tau\gamma}{K}$ in (14) and setting $\tilde{\sigma}=\sigma/\sqrt{K}$, we obtain (16).
A-C Proof of Theorem 13
Our proof relies on the following fact about the mixture of uniform distributions.
Lemma 15.
If $Z\,|\,\{\tilde{U}=u\}\sim\mathrm{Unif}((-bu,bu))$ and $\tilde{U}\sim\mathrm{Gamma}(2,1)$, then $Z\sim\mathrm{Lap}(0,b)$.
By setting $\tilde{U}_{t+\tau,j}^{k}\sim g=\mathrm{Gamma}(2,1)$ for every $k$ and $\tilde{Z}_{t+\tau,j}^{k}\;|\;\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}((-bu,bu))$ for $k$ and $j$, Lemma 11 and Lemma 15 imply that the resulting mechanism is Laplace.
The remaining proof is similar to that of Theorem 12, and the details are omitted. Note that the $\ell_{1}$-sensitivity $\Delta_{1}$ is at most $2\tau\gamma$.